# ILCL: Inverse Logic-Constraint Learning from Temporally Constrained Demonstrations

### v1

@@ Abstract
We aim to solve the problem of temporal-constraint learning from demonstrations to reproduce demonstration-like logic-constrained behaviors. Learning logic constraints is challenging due to the combinatorially large space of possible specifications and the ill-posed nature of non-Markovian constraints. To figure it out, we introduce a novel temporal-constraint learning method, which we call inverse logic-constraint learning (ILCL). Our method frames ICL as a two-player zero-sum game between 1) a genetic algorithm-based temporal-logic mining (GA-TL-Mining) and 2) logic-constrained reinforcement learning (Logic-CRL). GA-TL-Mining efficiently constructs syntax trees for parameterized truncated linear temporal logic (TLTL) without predefined templates. Subsequently, Logic-CRL finds a policy that maximizes task rewards under the constructed TLTL constraints via a novel constraint redistribution scheme. Our evaluations show ILCL outperforms state-of-the-art baselines in learning and transferring TL constraints on four temporally constrained tasks. We also demonstrate successful transfer to real-world peg-in-shallow-hole tasks.
@@ I Introduction
Robot policy learning often involves constraints that account for user, task, and environmental restrictions, as shown in Fig. 1. To autonomously adopt these constraints, inverse constraint learning (ICL) methods recover constraints from demonstration [1]. While ICL typically targets global constraints, most tasks require spatial or temporal restrictions. For example, mobile robots must wait until the traffic sign turns green. To express such structured and interpretable restrictions, researchers often adopt temporal logic (TL), such as linear temporal logic (LTL) [2] and signal temporal logic (STL) [3].
@@
In this context, we aim to solve the problem of TL-constraint learning from demonstrations to reproduce demonstration-like constraint-satisfying behaviors. However, TL-constraint learning is challenging due to the combinatorially large space of logic specifications in addition to the non-differentiable representation following strict syntactic rules. For efficient searches, researchers often rely on predefined logic templates [4] or restrict operator sets [5]. However, the limited expressivity restricts applications in tasks. Therefore, a desired approach requires to derive free-form logic constraints.
@@
Figure 1: An exemplar peg-in-shallow-hole task, which requires maintaining a peg-hole contact while inserting the peg using a parallel jaw gripper. By learning temporal logic constraints from demonstrations, our method ILCL successfully transfers demonstration-like constrained insertion behavior on a unseen tilted hole environment.
@@
Another challenge arises from the ill-posed nature of non-Markovian constraints. Given this ill-posedness, conventional approaches often focus on finding constraints without simultaneously considering the reward maximization in policy learning. This reduces the chance of finding optimal and transferable constraints from experts. Further, the non-Markovian property requires evaluating the history of state-action sequences, which yields sparse feedback during policy learning. These combined challenges complicate the integration of TL with conventional ICL frameworks [1] via inverse and forward constraint reinforcement learning (CRL) processes.
@@
We propose inverse logic-constraint learning (ILCL), a free-form transferable TL-constraint learning algorithm. Our method involves constraint learning as a two-player zero-sum game, similar to [6], between 1) a genetic algorithm-based temporal-logic mining (GA-TL-Mining) and 2) logic-constrained reinforcement learning (Logic-CRL). GA-TL-Mining efficiently constructs a syntax tree for parameterized truncated linear temporal logic (TLTL) [7] without predefined templates. Subsequently, Logic-CRL finds a policy that maximizes task rewards while adhering to the constructed TLTL constraints, measuring the degree of logic satisfaction across a trajectory using a robustness function. To handle the history dependency inherent in TL constraints in CRL, we introduce a constraint redistribution method, analogous to reward redistribution approaches [8, 9], converting trajectory-based constraints into state-action constraints. To the best of our knowledge, we provide the first CRL approach that generalizes to arbitrary TL constraints.
@@
We perform a statistical evaluation of ILCL against state-of-the-art ICL baselines. Across four simulated navigation and manipulation benchmark environments, ILCL outperforms baselines in learning diverse forms of TL constraints, achieving the lowest constraint-violation scores while yielding high task reward returns comparable to expert behaviors. We also demonstrate the applicability of the learned constraints by successfully transferring a constraint to a real-world peg-in-shallow-hole task, thereby demonstrating the robustness and scalability of TL constraints.
@@ II Related Work
Constraint learning and mining. ICL methods recover constraints from demonstrations. Early methods seek to learn numerical constraints in a discretized state space [10], while, with the advancement of neural representations, recent methods model constraints on continuous state or state-action spaces [11, 12]. To enhance model scalability, researchers introduce multi-task [6] and multi-modal [13] approaches. However, numerical representations are often hard to interpret and transfer to new settings. Instead, we use logical representations for robust generalization.
@@
Logic constraint learning is a variant of ICL. Early methods identify logic formulas adopting inductive logic programming [14], one-class decision tree [15], or data-driven rule induction [16]. However, these approaches often limit the outcomes to simple global propositions (e.g., ${}^{\forall}(x<3)$) or timestep-specific constraints, thereby restricting generalizability. To enhance temporal representation, researchers introduce STL learning, but the algorithms still face the same limitation of relying on predefined logic forms [4] or limited temporal expressions [5]. In contrast, our method employs specification mining without relying on pre-defined structures.
@@
Logic specification mining refers to the process of inferring potential system properties from observations (e.g., STL mining [17]). For effective inference in the combinatorially large specification space, researchers often fix logic structures  [18], simplify forms [19], discretize the search space [20], or restrict the search depth [21]. Alternatively, researchers enable tree-based genetic algorithms to explore the space of logic syntax trees exchanging subtrees without restrictions [22]. We extend the algorithm to compose a TL constraint with additional temporal operators, e.g., $\mathcal{R},\bigcirc$).
@@
Constrained Reinforcement Learning. CRL, also known as safety RL, aims to find a reward-maximizing policy while satisfying safety constraints. Garcia and Fernandez classify approaches into two categories: 1) modification of the exploration process and 2) the modification of the optimization criterion [23]. The first category typically restrict actions or policies within a trust region [24]. While these connect to formal verification, the temporal constraints we use make it hard to define the trust space for action sequences.
@@
On the other hand, the second category often employs Lagrangian relaxation techniques to convert the constrained optimization in CRL into a typical RL problem with unconstrained objectives (i.e., rewards) [25]. However, the long-term dependency in temporal logics may result in reward delays. To address this, researchers introduce neural networks to estimate trajectory returns [8] or redistribute rewards across time steps, smoothing the reward signals [9]. We extend this reward redistribution to TL-based CRL evaluating constraint violations at the end of episodes.
@@
In addition, TL-based CRL approaches often prioritize logic satisfaction, neglecting the primary task objective within environments. To mitigate this, approaches weigh logic-based rewards [26], apply logic fragments only [27], or discretize the environment into problem-specific state spaces [28]. These strategies limit their generalizability across novel environments and the adoption of complex TL specifications. In contrast, our method operates on arbitrary TL in a model-free manner.
@@ III-A Constraint Markov Decision Process (CMDP)
A CMDP is a tuple $(\mathcal{S},\mathcal{A},P,R,C)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{+}$ is the transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, and $C:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the constraint-cost function. A policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{+}$ defines a probability density function to determine the likelihood of selecting an action given a state. Then, the objective of CMDP is to find a policy $\pi^{*}$ that generates a state-action sequence $\xi=({\bf s}_{0},{\bf a}_{0},...{\bf s}_{T},{\bf a}_{T})$ with a finite horizon $T$ maximizing the expected return while the expectation of cumulative constraint cost bounded to a budget $d$: $\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\xi\sim\pi}[R({\bf s}_{t},{\bf a}_{t})]$ subject to $\mathbb{E}_{\xi\sim\pi}[\sum^{T}_{t=0}C({\bf s}_{t},{\bf a}_{t})]\leq d$. In this work, we use $R(\xi)=\sum^{T}_{t=0}R({\bf s}_{t},{\bf a}_{t})$ and $C(\xi)=\sum^{T}_{t=0}C({\bf s}_{t},{\bf a}_{t})$ to denote the cumulative rewards and constraint costs, respectively.
@@
Figure 2: Illustration of ILCL that learns TLTL constraint from demonstration to generalize to novel scenario. ILCL consists of two parts: GA-TL-Mining (Middle top) which generates a TLTL constraint that distinguishes demonstration trajectories from generated trajectories through a genetic algorithm on the TL syntax tree; and Logic-CRL, which optimizes the policy for task rewards under the generated TL constraint using SAC-Lag [29]. Once ILCL identifies a TL constraint, Logic-CRL derives a new policy applicable to novel scenarios.
@@ III-B Truncated Linear Temporal Logic (TLTL)
We employ TLTL [7], a predicate temporal logic that extends LTL [2] and LTL over finite traces ($\text{LTL}_{f}$ [30]). Each TLTL formula consists of atomic predicates (AP) of the form $a\cdot\mathbf{s}^{(i)}<b$, where $\mathit{a}\in\{-1,1\}$, $b\in\mathbb{R}$ is a constant, and $\mathbf{s}^{(i)}$ refers to the $i$-th element of the state vector ${\bf s}$. The syntax of TLTL formulas with an AP $\mu$ is defined as follows:
$|\displaystyle\phi::=\top\,|\,\bot$$\,|\,\mu\,|\,\neg\mu\,|\,\phi_{1}\land\phi_{2}\,|\,\phi_{1}\lor\phi_{2}\,|\,\bigcirc\phi\,|\,$
@@
To incorporate learnable thresholds into TLTL, we introduce a parametric variant, we call pTLTL, inspired by parametric STL (pSTL) [31]. The parameterization replaces the constant $b$ in each AP with a parameter $\theta$, yielding a parametric AP (pAP) of the form $a\cdot{\bf s}^{(i)}<\theta$. For example, we represent a TLTL formula $\phi$ as a pTLTL formula $\phi_{\boldsymbol{\theta}}$:
[…]
where a parameter vector $\boldsymbol{\theta}=(\boldsymbol{\theta}^{(0)},\boldsymbol{\theta}^{(1)})\in\mathbb{R}^{2}$. In this work, we use the learnable pTLTL; for simplicity, we use TLTL $\phi$ and pTLTL $\phi_{\boldsymbol{\theta}}$ interchangeably and denote them as $\phi$.
Then, the Boolean semantics of a formula $\phi$ defines a state sequence ${\bf s}_{t:t+\Delta}$ satisfies $\phi$ at time $t$ if and only if $\rho({\bf s}_{t:t+\Delta},\phi,t)>0$, where $\rho$ is the robustness function [7]. This function provides quantitative semantics, assigning large positive values for strong satisfaction and large negative values for significant violation. For simplicity, we denote robustness as $\rho(\xi,\phi)$, omitting the time index.
@@ IV-A Problem formulation
The objective of ILCL is to identify a constraint $C_{\phi}$, represented by a TLTL formula $\phi$, from expert trajectories $\Xi_{E}$. Then, we formulate the identification as a constrained min-max optimization that finds an optimal constraint $C_{\phi}^{*}$ maximizing the expected cumulative rewards of the the inferred constrained policy $\pi$, while minimizing its expectation difference from the expert policy $\pi_{E}$ demonstrated in $\Xi_{E}$:
$C_{\phi}^{*}=$ $\arg\min_{C_{\phi}}\max_{\pi}$ $\mathbb{E}_{\xi\sim\pi}[R(\xi)]-\mathbb{E}_{\xi\sim\pi_{E}}[R(\xi)]$
s.t. $\mathbb{E}_{\xi\sim\pi}[C_{\phi}(\xi)]=0$ (3)
$\mathbb{E}_{\xi\sim\pi_{E}}[C_{\phi}(\xi)]=0,$
where $C_{\phi}(\xi)=\mathds{1}[\rho(\xi,\phi)<0]$. However, this problem is ill-posed and also intractable due to the computation of $\pi$ for every candidate $C_{\phi}$.
Alternatively, we reformulate Eq. (IV-A) as a two-player zero-sum game, inspired by inverse state-action constraint learning in [6]. This approach is to mitigate degenerate conservative solutions in ICL. Our key distinction from [6] is the use of a logic-based constraint $C_{\phi}$ defined over state-action sequences. This formulation alternates between constraint and policy optimization players as shown in Fig. 2. In each $k$-th iteration, the constraint optimization (i.e., GA-TL-Mining) identifies a constraint $C_{\phi,k}$ that maximally penalizes the previously learned policies $\pi_{1},\dots,\pi_{k}$ relative to the expert policy $\pi_{E}$:
$C_{\phi,k}=$ $\mathop{\arg\max}\limits_{C_{\phi}}\;\sum_{i\leq k}\mathbb{E}_{\xi\sim\pi_{i}}\left[C_{\phi}(\xi)\right]$ (4)
[…]
The policy optimization (i.e., Logic-CRL) then finds a new optimal policy $\pi_{k+1}$ that satisfies the identified constraint $C_{\phi,k}$:
$\pi_{k+1}=$ $\mathop{\arg\max}\limits_{\pi}\;\mathbb{E}_{\xi\sim\pi}\left[R(\xi)\right]\;\textrm{s.t.}\;\mathbb{E}_{\xi\sim\pi}[C_{\phi,k}(\xi)]=0.$ (5)
These adversarial alternations progressively generate better-fitting constraints, narrowing the gap to the ground truth. The game’s convergence details are available in [6]. The identified constraint then induces a policy that minimally violates the constraint while gaining higher rewards than the expert policy. Below, we describe each player in detail.

### v2

@@ Abstract
We aim to solve the problem of temporal-constraint learning from demonstrations to reproduce demonstration-like logic-constrained behaviors. Learning logic constraints is challenging due to the combinatorially large space of possible specifications and the ill-posed nature of non-Markovian constraints. To this end, we introduce inverse logic-constraint learning (ILCL), a novel temporal-constraint learning method formulated as a two-player zero-sum game between 1) a genetic algorithm-based temporal-logic mining (GA-TL-Mining) and 2) logic-constrained reinforcement learning (Logic-CRL). GA-TL-Mining efficiently constructs syntax trees for parameterized truncated linear temporal logic (TLTL) without predefined templates. Subsequently, Logic-CRL finds a policy that maximizes task rewards under the constructed TLTL constraints via a novel constraint redistribution scheme. Our evaluations show ILCL outperforms state-of-the-art baselines in learning and transferring TL constraints on four temporally constrained tasks. We also demonstrate successful transfer to real-world peg-in-shallow-hole tasks.
@@ I Introduction
Robot policy learning often involves constraints that account for user, task, and environmental restrictions, as shown in Fig. 1. To autonomously adopt constraints, inverse constraint learning (ICL) methods recover constraints from demonstration [liu2025a]. While ICL typically targets global constraints, most tasks require spatial or temporal restrictions. For example, mobile robots must wait until the traffic sign turns green. To express such structured and interpretable restrictions, researchers often adopt temporal logic (TL), such as linear temporal logic (LTL) [ltlpnueli1977temporal] and signal temporal logic (STL) [stlmaler2004monitoring].
@@
In this context, we aim to solve the problem of TL-constraint learning from demonstrations to reproduce demonstration-like constraint-satisfying behaviors. However, TL-constraint learning is challenging due to the combinatorially large space of logic specifications in addition to the non-differentiable representation following strict syntactic rules. For efficient searches, researchers often rely on predefined logic templates [yifru2024concurrent] or restrict operator sets [liu2024interpretable]. However, the limited expressivity restricts applications in tasks. Therefore, a desired approach requires to derive free-form logic constraints.
@@
Figure 1: An exemplar peg-in-shallow-hole task, where one corner of the peg must maintain contact with the hole during insertion using a parallel jaw gripper. By learning temporal logic constraints from demonstrations, our method successfully transfers the constrained insertion behavior to an unseen, tilted hole environment.
@@
Another challenge arises from the ill-posed nature of non-Markovian constraints. Given this ill-posedness, conventional approaches often focus on finding constraints without simultaneously considering the reward maximization in policy learning. This reduces the chance of finding optimal and transferable constraints from experts. Further, the non-Markovian property requires evaluating the history of state-action sequences, which yields sparse feedback during policy learning. These combined challenges complicate the integration of TL with conventional ICL frameworks [liu2025a] via inverse and forward constraint reinforcement learning (CRL) processes.
@@
We propose inverse logic-constraint learning (ILCL), a free-form transferable TL-constraint learning algorithm. Our method involves constraint learning as a two-player zero-sum game [kim2024learning], between 1) a genetic algorithm-based temporal-logic mining (GA-TL-Mining) and 2) logic-constrained reinforcement learning (Logic-CRL). GA-TL-Mining constructs a syntax tree for parameterized truncated linear temporal logic (TLTL) [tltlli2017reinforcement] without predefined templates. Subsequently, Logic-CRL finds a policy that maximizes task rewards while adhering to the constructed TLTL constraints, measuring the degree of logic satisfaction across a trajectory using a robustness function. To handle the history dependency inherent in TL constraints in CRL, we introduce a constraint redistribution method, analogous to reward redistribution approaches [arjona2019rudder, gangwani2020learning], converting trajectory constraints into state-action constraints. To the best of our knowledge, we provide the first CRL approach that generalizes to arbitrary TL constraints.
@@
We evaluate ILCL against state-of-the-art ICL baselines in four simulated navigation and manipulation benchmark environments. ILCL outperforms baselines in learning diverse forms of TL constraints, achieving the lowest constraint-violation scores while yielding high task reward returns comparable to expert behaviors. We also demonstrate the applicability of the learned constraints by successfully transferring a constraint to a real-world peg-in-shallow-hole task, thereby demonstrating the robustness and scalability of TL constraints.
@@ II Related Work
Constraint learning and mining. ICL recovers constraints from demonstrations. Early work learns numerical constraints in discretized states [scobeemaximum], while, with neural representations, recent methods model constraints on continuous state or state-action spaces [malik2021inverse, jang2023inverse]. To scale, researchers introduce multi-task [kim2024learning] and multi-modal [qiao2024multi] approaches. However, numerical representations are often hard to interpret and transfer. Instead, we use logical representations for robust generalization.
@@
Logic constraint learning is a variant of ICL. Early methods identify logic formulas adopting inductive logic programming [baert2023learning], one-class decision tree [baert2024learning], or data-driven rule induction [kusters2022differentiable]. However, these approaches yield simple global propositions (e.g., $^{\forall}(x<3)$) or timestep-specific constraints, thereby restricting generalizability. To enhance temporal representation, researchers introduce TL in imitation learning, but algorithms still face the same limitation of relying on predefined logic forms [yifru2024concurrent], [wang2023temporal] or limited temporal expressions [liu2024interpretable]. In contrast, our method performs specification mining without relying on pre-defined templates.
@@
Logic specification mining infers candidate system properties from observations (e.g., STL mining [bartocci2022survey]). To handle the combinatorially large specification space, researchers often fix logic structures [jha2017telex], discretize the search space [vaidyanathan2017grid], or restrict the search depth [mohammadinejad2020interpretable]. Alternatively, researchers enable tree-based genetic algorithms to explore the space of logic syntax trees without restrictions [nenzi2018robust]. We extend the algorithm to compose a TL constraint with additional temporal operators.
@@
Constrained Reinforcement Learning. CRL, also known as safety RL, aims to find a reward-maximizing policy while satisfying safety constraints. Garcia and Fernandez classify approaches into two categories: 1) modification of the exploration process and 2) the modification of the optimization criterion [garcia2015comprehensive]. The first category typically restricts actions or policies within a trust region [achiam2017constrained]. Although these connect to formal verification, the temporal constraints we use make it hard to define the trust space for action sequences.
@@
On the other hand, the second category often employs Lagrangian relaxation to convert the constrained optimization in CRL into a typical RL problem with unconstrained objectives (i.e., rewards) [tessler2018reward]. However, the long-term dependency on TL may result in reward delays. To address this, researchers introduce neural networks to estimate trajectory returns [arjona2019rudder] or redistribute rewards across time steps, smoothing the reward signals [gangwani2020learning]. We extend this redistribution to TL-based CRL, evaluating constraint violations at the end of episodes.
@@
In addition, TL-based CRL approaches often prioritize logic satisfaction, neglecting the primary task objective within environments. To mitigate this, approaches weigh logic-based rewards [Shah2024ltlconstrained] or apply logic fragments only [Wen2015CorrectbysynthesisRL]. These strategies limit their generalizability across novel environments and the adoption of complex TL specifications. In contrast, our method operates on arbitrary TL in a model-free manner.
@@ III-A Constraint Markov Decision Process (CMDP)
A CMDP is a tuple $(\mathcal{S},\mathcal{A},P,R,C)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{+}$ is the transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, and $C:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the constraint-cost function. A policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{+}$ defines a probability density function to determine the likelihood of selecting an action given a state. Then, the objective of CMDP is to find a policy $\pi^{*}$ that generates a state-action sequence $\xi=({\bf s}_{0},{\bf a}_{0},...,{\bf s}_{T},{\bf a}_{T})$ with a finite horizon $T$ maximizing the expected return while the expectation of cumulative constraint cost bounded to a budget $d$: $\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\xi\sim\pi}[\sum^{T}_{t=0}R({\bf s}_{t},{\bf a}_{t})]$ subject to $\mathbb{E}_{\xi\sim\pi}[\sum^{T}_{t=0}C({\bf s}_{t},{\bf a}_{t})]\leq d$. In this work, we use $R(\xi)=\sum^{T}_{t=0}R({\bf s}_{t},{\bf a}_{t})$ and $C(\xi)=\sum^{T}_{t=0}C({\bf s}_{t},{\bf a}_{t})$ to denote the cumulative rewards and constraint costs, respectively.
@@
Figure 2: Illustration of ILCL that learns TLTL constraint from demonstration to generalize to novel scenario. ILCL consists of two parts: GA-TL-Mining (Middle top) which generates a TLTL constraint that distinguishes demonstration trajectories from generated trajectories through a genetic algorithm on the TL syntax tree; and Logic-CRL, which optimizes the policy for task rewards under the generated TL constraint using SAC-Lag [ha2021learning]. Once ILCL identifies a TL constraint, Logic-CRL derives a new policy applicable to novel scenarios.
@@ III-B Truncated Linear Temporal Logic (TLTL)
We employ TLTL [tltlli2017reinforcement], a predicate temporal logic that extends LTL [ltlpnueli1977temporal] and LTL over finite traces ($\text{LTL}_{f}$ [de2013linear]). Each TLTL formula consists of atomic predicates (AP) of the form $a\cdot\mathbf{s}^{(i)}<b$, where $a\in\{-1,1\}$, $b\in\mathbb{R}$ is a constant, and $\mathbf{s}^{(i)}$ refers to the $i$-th element of the state vector $\bf s$}. The syntax of TLTL formulas with an AP $\mu$ is defined as follows:
$\phi::=\top\,|\,\bot$ $\,|\,\mu\,|\,\neg\mu\,|\,\phi_{1}\land\phi_{2}\,|\,\phi_{1}\lor\phi_{2}\,|\,\bigcirc\phi\,|\,$
@@
To incorporate learnable thresholds into TLTL, we introduce a parametric variant, we call pTLTL, inspired by parametric STL (pSTL) [asarin2012parametric]. The parameterization replaces the constant $b$ in each AP with a parameter $\theta$, yielding a parametric AP (pAP) of the form $a\cdot{\bf s}^{(i)}<\theta$. For example, we represent a TLTL formula $\phi$ as a pTLTL formula $\phi_{\bm{\theta}}$:
$\phi$ $=({\bf s}^{(0)}<1)\,\mathcal{U}\,({\bf s}^{(1)}>2)$ $\dots\quad\,\,\,\text{(TLTL)},$
$\phi_{\bm{\theta}}$ $=({\bf s}^{(0)}<\bm{\theta}^{(0)})\,\mathcal{U}\,({\bf s}^{(1)}>\bm{\theta}^{(1)})$ $\dots\quad\text{(pTLTL)},$ (2)
where a parameter vector $\bm{\theta}=(\bm{\theta}^{(0)},\bm{\theta}^{(1)})\in\mathbb{R}^{2}$. In this work, we use the learnable pTLTL; for simplicity, we use TLTL $\phi$ and pTLTL $\phi_{\bm{\theta}}$ interchangeably and denote them as $\phi$.
Then, the Boolean semantics of a formula $\phi$ defines a state sequence ${\bf s}_{t:t+\Delta}$ satisfies $\phi$ at time $t$ if and only if $\rho({\bf s}_{t:t+\Delta},\phi,t)>0$, where $\rho$ is the robustness function [tltlli2017reinforcement]. This function provides quantitative semantics, assigning large positive values for strong satisfaction and large negative values for significant violation. For simplicity, we denote robustness as $\rho(\xi,\phi)$, omitting the time index.
@@ IV-A Problem formulation
The objective of ILCL is to identify a constraint $C_{\phi}$, represented by a TLTL formula $\phi$, from the expert trajectories $\Xi_{E}$ that exhibit zero violations. Then, we formulate the identification as a constrained min-max optimization that finds an optimal constraint $C_{\phi}^{*}$ maximizing the expected cumulative rewards of the inferred constrained policy $\pi$, while minimizing its expectation difference from the expert policy $\pi_{E}$:
$C_{\phi}^{*}=$ $\arg\min_{C_{\phi}}\max_{\pi}$ $\mathbb{E}_{\xi\sim\pi}[R(\xi)]-\mathbb{E}_{\xi\sim\pi_{E}}[R(\xi)]$ (3)
s.t. $\mathbb{E}_{\xi\sim\pi}[C_{\phi}(\xi)]=\mathbb{E}_{\xi\sim\pi_{E}}[C_{\phi}(\xi)]=0,$
where $C_{\phi}(\xi)\!\!=\!\!\mbox{$\mathbf{1}$}[\rho(\xi,\phi)\!\!<\!\!0]$}. However, this problem is intractable due to the computation of $\pi$ for every candidate $C_{\phi}$.
Alternatively, we reformulate Eq. (3) as a two-player zero-sum game, inspired by inverse state-action constraint learning in [kim2024learning]. Our key distinction from [kim2024learning] is the use of a logic-based constraint $C_{\phi}$ defined over state-action sequences. This formulation alternates between constraint and policy optimization players as shown in Fig. 2. In each $k$-th iteration, the constraint optimization (i.e., GA-TL-Mining) identifies a constraint $C_{\phi,k}$ that maximally penalizes the previously learned policies $\pi_{1},\dots,\pi_{k}$ relative to the expert policy $\pi_{E}$:
$C_{\phi,k}\hskip-2.0pt=$ $\;\underset{C_{\phi}}{\text{argmax}}\hskip-1.0pt\underset{i\leq k}{{\Sigma}}\mathbb{E}_{\xi\sim\pi_{i}}[C_{\phi}(\xi)],\,\textrm{s.t.}\,\,\mathbb{E}_{\xi\sim\pi_{E}}[C_{\phi}(\xi)]=0.$ (4)
The policy optimization (i.e., Logic-CRL) then finds a new optimal policy $\pi_{k+1}$ that satisfies the constraint $C_{\phi,k}$:
$\pi_{k+1}\hskip-2.0pt=\;$ $\underset{\pi}{\text{argmax}}\;\mathbb{E}_{\xi\sim\pi}\left[R(\xi)\right],\;\textrm{s.t.}\;\mathbb{E}_{\xi\sim\pi}[C_{\phi,k}(\xi)]=0.$ (5)
These alternations progressively find a constraint that separates the expert policy from discovered high reward policies than the expert. This aligns with the ground-truth constraint, as the expert policy obtains the highest reward among policies that satisfy the constraint. As a result, the identified constraint induces a policy that minimally violates the constraint while gaining higher rewards than the expert policy.